Answer questions 7, 15, 17

Answer questions 7, 15, 17

media%2F206%2F206463f5-947e-49d4-b6d7-28
Answer #1

7. given

n = 500 and p = 0.4

this gives q = 1-p = 0.6

so now,

mean \mu = np = 200 and \sigma = \sqrt{npq} = \sqrt{500*0.4*0.6}=10.95